OpenStudy (anonymous):

PART1 Simplify completely: the quantity 5 times x to the third power plus 10 times x to the 2nd power plus 15 times x all over 5 times x a.x^2 + 10x^2 + 15x b.2x + 6 c.x^2 + 2x + 3 d.x^4 + 2x^3 + 3x^2 PART2 Simplify completely: the quantity 3 times x to the 4th power plus 5 times x to the 2nd power plus 2 times x all over x a.the quantity 3 times x to the third power plus 5 times x plus 2 all over x b.3x^3 + 5x + 2 c.3x^2 + 5x^2 + 2x d.3x^5 + 5x^3 + 2x^2

OpenStudy (anonymous):

P1: (5x^3+10x^2+15x)/5x =x^2+5x+3

OpenStudy (anonymous):

2x**

OpenStudy (campbell_st):

the rule is: when dividing, subtract the powers of the same pronumeral a^n/a^m = a^(n-m) use the rule to learn what to do

OpenStudy (anonymous):

so P1: C

jhonyy9 (jhonyy9):

mattt9 the second resulted term is 2x

OpenStudy (anonymous):

i corrected that already. sorry.

OpenStudy (anonymous):

did anyone figure out part 2?